OpenStudy (anonymous):

A line passes through the point (0, 3) and has a slope of 5/4 . What is the equation of the line?

OpenStudy (solomonzelman):

you know that it has slope of 5/4, say y=5/4x+b, and plug in your point, (0, 3) to solve for b (3)=5/4(0)+b 3=b so y=5/4x+3

OpenStudy (anonymous):

A. 5x – 4y = –12 B. 5x + 4y = 12 C. 5x – 4y = –3 D.5x + 4y = 3

OpenStudy (solomonzelman):

y=5/4x+3 times 4 4y=5x+12 or -5x+4y=12 which is same as 5x-4y=-12 it is A

OpenStudy (anonymous):

thank you